B. Interpreting chemical equations using a balanced chemical equations: Solve for the required value:

1. How many moles of CO2 can be produced from 7.75 mole C2H5OH (Balance the equation first.) C2H5OH + O2______ CO2 + H2O


2. Ethane gas, C2H6, burns in air to form carbon dioxide and water. C2H6 + O2 ______ CO2 + H2O a. How many moles of O2 are needed for the complete combustion of 15.0 mole of ethane? b. How many grams of CO2 are produced for each 8.0 g H2O produced?




1
Expert's answer
2020-11-02T08:46:56-0500

C2H5OH + 3O2 = 2CO2 + 3H2O

n (CO2) = 7,75 * 2/1 = 15,5 mol

Anser: 15,5 mol


2C2H6+ 7O2= 4CO2 + 6H2O

n (O2) = 15 * 7/2 = 52,5 mol

Anser:52,5 mol


n (CO2) = 30 mol

m(CO2) ="n*M=" 1320 g

n (H2O) = 45 mol

m (H2O) = 810 g

810 g H2O – 1320 g CO2

8 g - X

Х = 8 * 1320/810 = 13 g CO2

Anser: 13 g